OpenStudy (anonymous):

Which of the following is the solution of –6x – 17 = 19?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Use the idea of balance to solve it :)

OpenStudy (anonymous):

-6x - 17 = 19 +17 +17 -6x = 36 x = -6

OpenStudy (skullpatrol):

@AlexaLeyva are you still there?

OpenStudy (anonymous):

@skullpatrol yeah

OpenStudy (anonymous):

Did you get my answer? @AlexaLeyva

OpenStudy (anonymous):

@darthb10 yeah

OpenStudy (anonymous):

Ok good :D. If you need anymore help just ask away.

OpenStudy (anonymous):

okay thank you(: